OpenStudy (anonymous):

calculate. - 1/4 - 9 =? is this not - 13/4 ? quiz saying wrong answer

OpenStudy (anonymous):

no it's -37/4

OpenStudy (anonymous):

can you explain please?

OpenStudy (anonymous):

dp -1/4 - 9/1 and that will give you -37/4

OpenStudy (anonymous):

9= 36/4. -1/4 - 36/4 = -37/4

OpenStudy (anonymous):

thanks guys